OpenStudy (anonymous):

Find the sum and choose the correct answer. 2n3 + 4n2 - 7 and -n3 + 8n - 9 -n3 - 4n2 - 8n - 16 n3 + 4n2 + 8n - 16 n3 + 4n2 + 8n + 16 3n3 + 4n2 + 8n - 2

OpenStudy (ajprincess):

Add the like terms

OpenStudy (anonymous):

\[(2n^3 + 4n^2 - 7) + (-n^3 + 8n - 9)\] Combine the like terms! 2n^3 + -1n^3 = 1n^3 = n^3 So as head start I'll tell you that it'll begin with n^3

OpenStudy (anonymous):

ow solve the rest by yourself and remember that 4n^2 and 8n cannot be combined!

OpenStudy (anonymous):

*now

OpenStudy (ajprincess):

2n3 + 4n2 - 7 +( -n3 + 8n - 9) =(2n^3-n^3)+4n^2+8n+(-7-9)

OpenStudy (ajprincess):

nw try.

OpenStudy (anonymous):

ok um ok im not sure

OpenStudy (anonymous):

i hate algebra

OpenStudy (anonymous):

I do too don't worry! I'm glad I already finished with it this year:) Ok, Your are going to keep the +4n^2 + 8n the way they are: n^3 + 4n^2 + 8n -7 + (-9) = -16 so your equation is: \[n^3 + 4n^2 + 8n - 16\]

OpenStudy (anonymous):

Which is option B :) I hope this helps! If you don't understand it let me know:)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

im getting it i think

OpenStudy (anonymous):

Soon you will understand it better...I took algebra 1 this year and hated it but after you get to more difficult lessons with Algebra this will seem easier:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you much